# MathMathgg

### v1

@@ IV Computational Methods
In this section, we introduce two methods used in this work, the Monte Carlo method and the sparse-matrix diagonalization method, to calculate the RGTME
@@
$\mathcal{J}=\frac{1}{\sqrt{2}}(J_{5,+}^{x}+iJ_{5,+}^{y}),$ (23)
where $J_{5,+}^{x/y}$ is the $x/y$-component of the charge-raising axial current. Since $\mathcal{J}$ is a rank-one spherical tensor carrying the magnetic quantum number $m_{\mathcal{J}}=1$, the Clebsch-Gordan coefficient $\langle J_{i}M_{i},11|J_{f}M_{f}\rangle$ appears in the denominator for the extraction of the reduced matrix element. For the processes studied in this work, we have $J_{i}=J_{f}=\frac{1}{2}$ for $\rightarrow^{3}$e $\beta$-decay and $J_{i}=0,J_{f}=1$ for $\rightarrow^{6}$i $\beta$-decay, and in our calculation, we set $M_{i}=-1/2$, $M_{f}=1/2$ for the former transition and $M_{i}=0$, $M_{f}=1$ for the latter.
For convenience in later analysis, we decompose $\mathcal{J}$ into different orders according to Fig. 1:
@@ IV.1 Monte Carlo Method
for the full Hamiltonian $H$ defined in Eq. (1), where $a_{t}$ is the temporal lattice spacing. To facilitate perturbative calculations that will be introduced below, we also define a non-perturbative transfer matrix
@@
corresponding to the non-perturbative Hamiltonian $H_{0}$ defined in Eq. (13).
@@
at large $t$. To evaluate it, auxiliary field transformations are applied to both the numerator and denominator, reducing the problem to a multi-dimensional integral that can be handled using Monte Carlo methods. A detailed discussion of this process is provided in Appendix C. However, as is discussed in Section II.2, a direct auxiliary field transformation on the full transfer matrix $M$ leads to severe sign problems, making the results highly noisy.
@@
with the variable $t$ suppressed for brevity. Here, $n$ denotes the order of expansion in $\Delta M$. The contributions from different orders of the current are written as LO, N2LO and N3LO, respectively. For the LO current, we retain the first two terms in the perturbative series and neglect higher-order corrections, namely,
@@
When designing the algorithm, we note that any $\beta$-decay process involves two different nuclei, making this essentially a two-channel problem. To control statistical errors, an appropriate probability distribution function is required to update the auxiliary fields during the Markov process. We leave relevant details in Appendix C.
@@ IV.2 Lanczos Eigenvector Method
For the $\rightarrow{{}^{3}}$e process, the dimension of the Hilbert space is not too large and a rigorous calculation is possible. Therefore, in addition to the Monte Carlo approach, we employ the Lanczos eigenvector technique as an alternative method Epelbaum et al. (2010b). Specifically, we directly diagonalize the full Hamiltonian $H$ to obtain the ground-state wave functions of 3H and 3He. These wave functions are then used as input to compute the RGTME following Eq. (22). We make the calculation on a box with its length $L\approx 20$ fm, which is sufficiently large to neglect finite-volume effects Elhatisari et al. (2024b).
The main advantage of the Lanczos method is its accuracy. Unlike Monte Carlo simulations, it is free from statistical errors and systematic errors from the perturbative method. Additionally, due to its precision, this method enables a precise determination of the three-body LECs $c_{D}$ and $c_{E}$ from 3H $\beta$-decay. Specifically, we fit them to the experimental binding energy and half-life of 3H, closely following the procedure in Refs. Gazit et al. (2009); Baroni et al. (2016b, 2018). The details are presented in the next section.
@@ V Results
V.1 3H $\beta$-decay
The details of 3H $\beta$-decay are discussed in this subsection. An essential step here is to convert the half life, $t_{1/2}$, into the RGTME. We make use of the formula Simpson (1987); Hardy et al. (1990); Schiavilla et al. (1998)
$(1+\delta_{R})t_{1/2}f_{V}=\frac{2\pi^{3}\mathrm{ln}2/[m_{e}^{5}V_{ud}^{2}G_{F}^{2}(1+\Delta_{R}^{V})]}{\mathrm{F}^{2}+(f_{A}/f_{V})g^{2}_{A}\mathrm{GT}^{2}},$ (35)
which expresses $t_{1/2}$ in terms of the Fermi matrix element
$\mathrm{F}=\sum_{n=1}^{3}\langle^{3}\mathrm{He}||\tau_{n,+}||^{3}\mathrm{H}\rangle$ (36)
and the RGTME. In Eq. (35), $m_{e}=0.511$ MeV is the electron mass, $V_{ud}=0.9737(3)$ is the up-down quark-mixing element of the Cabibbo-Kobayashi-Maskawa matrix Hardy and Towner (2020a), $G_{F}=1.1664\times 10^{-5}$ GeV-1 is the Fermi constant Tishchenko et al. (2013), $f_{V}=2.8355\times 10^{-6}$ and $f_{A}=2.8505\times 10^{-6}$ are Fermi functions Simpson (1987), $\delta_{R}=1.9\%$ is the outer radiative correction of 3H $\beta$-decay Raman et al. (1978), and $\Delta_{R}^{V}=0.02467(22)$ is the inner radiative correction Seng et al. (2018). Our lattice result of the Fermi matrix element is
$\mathrm{F}=0.9996,$ (37)
combined with the measured comparative half-life, $(1+\delta_{R})t_{1/2}f_{V}=1134.6(31)$Simpson (1987), determines the value of the RGTME:
$\mathrm{GT}_{\mathrm{exp}}=1.6473(41).$ (38)
The subscript “$\mathrm{exp}$” means that the result is derived from experimental data, although the lattice prediction of the Fermi matrix element is used as an input. We incorporate the errors in $g_{A},V_{ud}$ and $\Delta_{R}^{V}$, therefore the uncertainty above is larger than the one in Ref. Elhatisari et al. (2024b).
Now we show the determination of $c_{D}$ and $c_{E}$ using the Lanczos eigenvector method. In Fig. 2, $c_{E}$ is plotted against $c_{D}$ by fitting to the experimental binding energy $E(^{3}$)=8.482)=-8.482) = - 8.482 MeV. Similar to previous studies, we observe a nearly linear relation between $c_{D}$ and $c_{E}$ Gazit et al. (2009); Baroni et al. (2016b, 2018). Then we calculate the lattice RGTME, $\mathrm{GT}_{\mathrm{lat}}$, and plot its ratio with respect to $\mathrm{GT}_{\mathrm{exp}}$ in Fig. 3, as a function of $c_{D}$. All the points in Fig. 3 share the same values of $(c_{D},c_{E})$ with the ones in Fig. 2. Again, GTth/GTexp is a nearly linear function of $c_{D}$, consistent with Ref. Baroni et al. (2018). The values of $c_{D}$ and $c_{E}$ are determined as
@@
LO NNLO N3LO(OPE) N3LO(CT)
LE 1.6747 $-0.0083$ $-0.0480$ 0.0289
MC 1.6652(121) $-0.0071(14)$0.0509(37) 0.0273(22)
Table 2: Contributions to RGTME from currents at different orders for $c_{D}=$1.01, $c_{E}$=0.81. The upper and lower lines are the results of the Lanczos eigenvector method(LE) and the Monte Carlo (MC) method, respectively. The value in the bracket is the statistical error.
@@
LO NNLO N3LO(OPE) N3LO(CT) Total-LO Total Exp Knecht et al. (2012)
@@
Table 3: Contributions to RGTME of 6He $\beta$-decay from currents at different orders. Columns labeled with LO, NNLO, N3LO(OPE) and N3LO(CT) correspond to the diagrams illustrated in panel (a),(b),(c) and (d) of Fig. 1, respectively. The full result is given in the column labeled “Total”, while results including only corrections beyond LO are listed under “Total-LO”. The value in the bracket is the statistical uncertainty. The experimental value in the last column is cited from Ref. Knecht et al. (2012).
To evaluate the contributions from currents at different orders , we calculate the corresponding RGTMEs and present the results in Table 2. It is evident that the LO current provides the dominant contribution, while corrections from higher-order currents remain below 5%. To validate the two computational approaches described in Sec. IV, we also performed Monte Carlo calculations for 3H $\beta$-decay. The results of both approaches, shown in the second and third rows of Table 2, exhibit good agreement.
@@
V.2 6He $\beta$-decay
We first present the energies and charge radii of 6He and 6Li in Table. 4. The energy is computed up to second-order perturbation theory Lu et al. (2022), leading to relatively large uncertainties. The charge radius, on the other hand, is calculated up to the first order. Overall, our results show satisfactory agreement with experimental data.
Nuclei($J^{P}$) 6He(0+) 6Li(1+)
$E$/MeV $R_{\mathrm{ch}}$/fm $E$/MeV $R_{\mathrm{ch}}$/fm
Lat -28.80(1.21) 1.988(32) -31.88(93) 2.534(18)
Exp -29.27 2.068(11) -31.99 2.589(39)
@@
Table 3 presents the lattice predictions for the RGTMEs of 6He $\beta$-decay, including the total and partial contributions from different orders of currents. Similar to 3H $\beta$-decay, the LO term remains dominant, while higher-order corrections contribute less than 5%. When compared with experiment, the LO result is slightly larger, and the negative correction from higher-order currents shifts the RGTME towards the correct direction and improves its agreement with the experimental data.
Since the contribution of higher-order currents are only included with the zeroth-order perturbation theory, as presented in Eq. (33), it is necessary to assess the impact of this truncation of the perturbative series. Our calculations indicate that the first-order perturbative correction to the LO current remains below 10% of the zeroth-order contribution. Assuming this is also the case for higher-order currents, the truncation would affect the total RGTME by about 1%, which is under good control and doesn’t influence our main conclusions.
VI Summary
In this work, we investigate nuclear $\beta$-decays using lattice quantum Monte Carlo simulations within the framework of the nuclear lattice effective field theory (NLEFT). The three-body low-energy constants (LECs) $c_{D}$ and $c_{E}$ are determined from the binding energy and the half-life of 3H. Our prediction for 6He $\beta$-decay shows reasonable agreement with experimental data and validates the method employed in this work.
To conclude, this work establishes a foundation for investigating $\beta$-decays in light and medium-mass nuclei with $A\geq 4$ using NLEFT. We anticipate that the recently developed N3LO interaction Li et al. (2018); Elhatisari et al. (2024a); Liu et al. will improve the predictions. Additionally, our approach opens the door to studying other nuclear electroweak processes of physical interest Engel and Menéndez (2017); Cirigliano et al. (2018a, b); Hardy and Towner (2020b); Cirigliano et al. (2024a, b); Gennari et al. (2025) within a unified lattice field theory framework, complementing lattice QCD studies at the quark and gluon level Tuo et al. (2019); Detmold and Murphy (2020); Davoudi et al. (2024); Feng et al. (2020); Ma et al. (2024b).
Acknowledgements.
We are grateful for helpful discussions on the nuclear interactions with Dean Lee, Ulf-G. Meißner and Shihang Shen. We appreciate Hermann Krebs for the choice of axial currents, Young-Ho Song for the design of Monte Carlo sampling method and Jiangming Yao for helpful comments. X.F. and T.W. were supported in part by NSFC of China under Grant No. 12125501, No. 12293060, No. 12293063, and No. 12141501, and National Key Research and Development Program of China under No. 2020YFA0406400. B.L. was supported by NSAF No.U2330401 and National Natural Science Foundation of China with Grant No.12275259.
@@ Appendix A Nucleon-nucleon Scattering
We use the spherical wall method to calculate nucleon-nucleon phase shifts on the lattice SuppNN1 ; SuppNN2 ; SuppNN3 . The fitting result of $n$-$p$ scattering for the full Hamiltonian $H$ (blue dash-dot line) and the non-perturbative Hamiltonian $H_{0}$ (black dashed line) is shown in Fig. S 1, with the center-of-mass momentum $P\leq 200$ MeV. We also show the phase shifts extracted from the Nijmegen partial wave analysis SuppNN4 represented by the red points. For $H_{0}$, it is only fitted to ${}^{1}S_{0}$ and ${}^{3}S_{1}$ channels, so reasonable agreement with the experimental data appears in these two channels. The results for other channels are predictions based on the $S$-waves fit, and the agreement is not very satisfactory. We could in principle make an improvement by increasing the value $\Lambda^{\prime}_{\pi}$ in Eq. (13) or adding more operators in $H_{0}$, but this would introduce additional sign problems. For $H$, it shows good agreement with the experiment in most channels. In several $D$-wave channels there are some deviations, which can be eliminated if operators beyond N2LO are included.
Figure S 1: $n$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ for $H$ and $H_{0}$. The red points are from Nijmegen partial wave analysis. The blue and black lines denote fitting results of the full Hamiltonian $H$ and the non-perturbative Hamiltonian $H_{0}$, respectively.
Figure S 2: $p$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ in ${}^{1}S_{0}$ channel. The black points with errors are from Nijmegen partial wave analysis. The gray line denotes the fit result.
For the isospin-breaking potential $V_{pp}$ and $V_{nn}$, we tune the value of $C_{pp}$ to reproduce the $p$-$p$ phase shifts in ${}^{1}S_{0}$ channel SuppNN4 . The fitting result is given in Fig. S 2. For $V_{nn}$, we fit it to the experimental value of the $n$-$n$ scattering length SuppNN6 and effective range SuppNN7 in ${}^{1}S_{0}$ channel:

### v2

@@ IV Computational Methods
In this section, we introduce two methods used in this work, i.e., the Monte Carlo method and the sparse-matrix diagonalization method, to calculate the RGTME following the convention of Edmonds Edmonds (1960),
@@
$\mathcal{J}=-\frac{1}{\sqrt{2}}(A_{x}+iA_{y}),$ (21)
[…]
We comment that the definition of the RGTME, Eq. (20), is followed throughout the work. The same convention has also been adopted to study 6He $\beta$ decay in Ref. Gnech et al. (2021); King et al. (2020), so our result of 6He’s RGTME can be directly compared to theirs. For 3H $\beta$ decay, however, different conventions have been adopted by different works. For example, our definition of the RGTME is larger than the Gamow-Teller matrix element defined in Ref. Baroni et al. (2016b, 2018); Gnech et al. (2021); Elhatisari et al. (2024b) by a factor of $\sqrt{2}$, but the same as the one defined in Ref. Li Muli et al. (2025). One should be aware of this difference when making a comparison.
For latter convenience, we decompose $\mathcal{J}$ into different orders according to Fig. 1,
@@ IV.1 Monte Carlo method
for the full Hamiltonian $H$ defined in Eq. (1), where $a_{t}$ is the temporal lattice spacing. To facilitate perturbative calculations that will be introduced below, we also define a nonperturbative transfer matrix
@@
corresponding to the nonperturbative Hamiltonian $H_{0}$ defined in Eq. (13).
@@
at large $t$. To evaluate it, auxiliary field transformations are applied to both the numerator and denominator, reducing the problem to a multidimensional integral that can be handled using Monte Carlo methods. A detailed discussion of this process is provided in Appendix C. However, as discussed in Sec. II.2, a direct auxiliary field transformation on the full transfer matrix $M$ leads to severe sign problems, making the results highly noisy.
@@
with the time variable $t$ suppressed for brevity. Here, $n$ denotes the order of expansion in $\Delta M$. The contributions from different orders of the current are written as LO, N2LO, and N3LO, respectively. For the LO current, we retain the first two terms in the perturbative series and neglect higher-order corrections, namely,
@@
When designing the algorithm, we note that any $\beta$-decay process involves two different nuclei, making this essentially a two-channel problem. To control statistical errors, an appropriate probability distribution function is required to update the auxiliary fields during the Markov process. We leave relevant details to Appendix C.
@@ IV.2 Lanczos eigenvector method
[…]
The main advantage of the Lanczos method is its accuracy. Unlike Monte Carlo simulations, it is free from statistical errors and systematic errors from the perturbative method. Additionally, due to its precision, this method enables a precise determination of the three-body LECs $c_{D}$ and $c_{E}$ from 3H $\beta$ decay. Specifically, we fit them to the experimental binding energy and half-life of 3H, closely following the procedure in Refs. Gazit et al. (2009); Baroni et al. (2016b, 2018); Elhatisari et al. (2024b). The details are presented in the next section.
@@ V Results
V.1 3H $\beta$ decay
The details of 3H $\beta$ decay are discussed in this subsection. An essential step here is to convert the half-life, $t_{1/2}$, into the RGTME. We make use of the formula Simpson (1987); Hardy et al. (1990); Schiavilla et al. (1998); Chou et al. (1993)
$(1+\delta_{R})t_{1/2}f_{V}=\frac{2\pi^{3}\mathrm{ln}2/[m_{e}^{5}V_{ud}^{2}G_{F}^{2}(1+\Delta_{R}^{V})]}{B(\mathrm{F})+B(\mathrm{GT})f_{A}/f_{V}},$ (33)
which expresses $t_{1/2}$ in terms of the Fermi transition strength
$B(\mathrm{F})=\left(\sum_{n=1}^{3}\langle^{3}\mathrm{He}|\tau_{n,+}|^{3}\mathrm{H}\rangle\right)^{2},$ (34)
and the Gamow-Teller transition strength $B(\mathrm{GT})$, quadratic in the RGTME Chou et al. (1993),
$B(\mathrm{GT})=\frac{1}{2}g_{A}^{2}\mathrm{GT}^{2}.$ (35)
In Eq. (33), $m_{e}=0.511$ MeV is the electron mass, $V_{ud}=0.9737(3)$ is the up-down quark-mixing element of the Cabibbo-Kobayashi-Maskawa matrix Hardy and Towner (2020), $G_{F}=1.1664\times 10^{-5}$ GeV-1 is the Fermi constant Tishchenko et al. (2013), $f_{V}=2.8355\times 10^{-6}$ and $f_{A}=2.8505\times 10^{-6}$ are Fermi functions Simpson (1987), $\delta_{R}=1.9\%$ is the outer radiative correction of 3H $\beta$ decay Raman et al. (1978), and $\Delta_{R}^{V}=0.02467(22)$ is the inner radiative correction Seng et al. (2018). Our lattice result of the Fermi transition strength is
$B(\mathrm{F})=0.9996^{2}.$ (36)
Note that because of vector current conservation, $B(\mathrm{F})$ is strictly equal to 1 in the isospin-symmetric limit. In our calculation, the isospin-breaking potential $V_{\mathrm{IB}}$ introduced in Eqs. (7), (8) and (9) generates a small shift, at the level of $10^{-4}$. Combining the Fermi transition strength Eq. (36) with the measured comparative half-life, $(1+\delta_{R})t_{1/2}f_{V}=1134.6(31)$Simpson (1987), the value of the RGTME is determined as
$\mathrm{GT}_{\mathrm{exp}}=2.3297(58).$ (37)
The subscript $\mathrm{exp}$” means that the result is derived from experimental data, although the lattice prediction of the Fermi transition strength is used as an input. Note that our definition of 3H’s RGTME is different from Ref. Gazit et al. (2009); Baroni et al. (2016b, 2018); Elhatisari et al. (2024b) by an overall constant $\sqrt{2}$, as has been emphasized in the last section.
Now we show the determination of $c_{D}$ and $c_{E}$ using the Lanczos eigenvector method. In Fig. 2, $c_{E}$ is plotted against $c_{D}$ by fitting to the experimental binding energy $E(^{3}$H)=8.482)=-8.482 MeV. Similar to previous studies, we observe a nearly linear relation between $c_{D}$ and $c_{E}$ Gazit et al. (2009); Baroni et al. (2016b, 2018). Then we calculate the lattice RGTME, $\mathrm{GT}_{\mathrm{lat}}$, and plot its ratio with respect to $\mathrm{GT}_{\mathrm{exp}}$ in Fig. 3, as a function of $c_{D}$. All the points in Fig. 3 share the same values of $(c_{D},c_{E})$ with the ones in Fig. 2. Again, GTth/GTexp is a nearly linear function of $c_{D}$, and the values of $c_{D}$ and $c_{E}$ are determined as
@@
Note that low-energy nucleon-deuteron ($n$-$d$) scattering provides an alternative process to help pin down $c_{D}$ and $c_{E}$ and to test the result above from 3H $\beta$ decay. The strategy has been employed by several works in the continuum  Epelbaum et al. (2002); Piarulli et al. (2018); Epelbaum et al. (2019), but it should be understood that a direct comparison of $c_{D}$ and $c_{E}$ cannot be performed due to different regularization schemes and regulators used in our work and in  Epelbaum et al. (2002); Piarulli et al. (2018); Epelbaum et al. (2019). In NLEFT, the adiabatic projection method  Rokash et al. (2015); Elhatisari (2019) was applied to investigating $n$-$d$ scattering Pine et al. (2013); Elhatisari et al. (2016b) using LO pionless EFT, with no discussion on the determination of three-body forces. Thus, a more systematic study of $n$-$d$ scattering on the lattice is required.
@@
LO N2LO N3LO(OPE) N3LO(CT)
LE 2.3684 -0.0117 -0.0679 0.0409
MC 2.355(17) -0.010(2) -0.070(1) 0.039(3)
Table 2: Contributions to RGTME from currents at different orders for $c_{D}=$1.01, $c_{E}$=0.81. The upper and lower lines are the results of the Lanczos eigenvector method (LE) and the Monte Carlo (MC) method, respectively. The value in parentheses is the statistical error.
Nuclei($J^{P}$) 6He(0+) 6Li(1+)
EE/MeV $R_{\mathrm{ch}}$/fm EE/MeV $R_{\mathrm{ch}}$/fm
Lattice -28.80(1.21) 1.988(32) -31.88(93) 2.534(18)
Expt. -29.27 2.068(11) -31.99 2.589(39)
@@
LO N2LO N3LO(OPE) N3LO(CT) Total-LO Total Expt. Knecht et al. (2012)
@@
Table 4: Contributions to RGTME of 6He $\beta$ decay from currents at different orders. Columns labeled with LO, N2LO, N3LO(OPE), and N3LO(CT) correspond to the diagrams illustrated in panels (a), (b), (c), and (d) of Fig. 1, respectively. The full result is given in the column labeled “Total,” while results including only corrections beyond LO are listed under “Total-LO.” The value in parenthese is the statistical uncertainty. The experimental value in the last column is from Ref. Knecht et al. (2012).
To evaluate the contributions from currents at different orders , we calculate the corresponding RGTMEs and present the results in Table 2. It is evident that the LO current provides the dominant contribution, while corrections from higher-order currents remain below 5%. To validate the two computational approaches described in Sec. IV, we also performed Monte Carlo calculations for 3H $\beta$ decay. The results of both approaches, shown in the second and third rows of Table 2, exhibit good agreement.
@@
V.2 6He $\beta$ decay
We first present the energies and charge radii of 6He and 6Li in Table 3. The energy is computed up to second-order perturbation theory Lu et al. (2022), leading to relatively large uncertainties. The charge radius, on the other hand, is calculated up to the first order. Overall, our results show satisfactory agreement with experimental data.
Next, we present the results for 6He $\beta$ decay. The dependence of the RGTME on the Euclidean evolution time $t$ is illustrated in Fig. 4, where contributions from currents at LO (black points) and beyond LO (blue points) are plotted separately in the upper and lower panels, respectively. The full result is represented by the red points in the upper panel. For the LO current, we observe a monotonic increase with $t$ and find a plateau beginning at around $t=0.3$ MeV-1. In contrast, the net contribution beyond LO exhibits a relatively weaker dependence on $t$. Therefore, we make a constant fit in the interval 0.3 $\leq t\leq$ 0.5 MeV-1 for the LO results and 0.15 $\leq t\leq$ 0.5 MeV-1 for the beyond-LO results. The fit result for the full RGTME is plotted as the red band in the upper panel, and is very close to its experimental value represented by the horizontal dashed purple line.
Table 4 presents the lattice predictions for the RGTMEs of 6He $\beta$-decay, including the total and partial contributions from different orders of currents. Similarly to 3H $\beta$ decay, the LO term remains dominant, while higher-order corrections contribute less than 5%. When compared with experiment, the LO result is slightly larger, and the negative correction from higher-order currents shifts the RGTME towards the correct direction and improves its agreement with the experimental data.
Since the contributions of higher-order currents are only included with the zeroth-order perturbation theory, as presented in Eq. (31), it is necessary to assess the impact of this truncation of the perturbative series. Our calculations indicate that the first-order perturbative correction to the LO current remains below 10% of the zeroth-order contribution. Assuming this is also the case for higher-order currents, the truncation would affect the total RGTME by about 1%, which is under good control and does not influence our main conclusions.
VI Summary
In this work, we investigate nuclear $\beta$ decays using lattice quantum Monte Carlo simulations within the framework of the nuclear lattice effective field theory (NLEFT). The three-body low-energy constants (LECs) $c_{D}$ and $c_{E}$ are determined from the binding energy and the half-life of 3H. Our prediction for 6He $\beta$ decay shows reasonable agreement with experimental data and validates the method employed in this work.
To conclude, this work establishes a foundation for investigating $\beta$ decays in light and medium-mass nuclei with $A\geq 4$ using NLEFT. We anticipate that the recently developed N3LO interaction Li et al. (2018); Elhatisari et al. (2024a); Liu et al. (2025) will improve the predictions. Additionally, our approach opens the door to studying other nuclear electroweak processes of physical interest Engel and Menéndez (2017); Cirigliano et al. (2018a, b); Hardy and Towner (2020); Cirigliano et al. (2024a, b); Gennari et al. (2025) within a unified lattice field theory framework, complementing lattice QCD studies at the quark and gluon level Tuo et al. (2019); Detmold and Murphy ; Davoudi et al. (2024); Feng et al. (2020); Ma et al. (2024b).
Acknowledgements.
We are grateful for helpful discussions on the nuclear interactions with Dean Lee, Ulf-G. Meißner, and Shihang Shen. We appreciate Hermann Krebs for the choice of axial currents, Young-Ho Song for the design of Monte Carlo sampling method, and Jiangming Yao for helpful comments. X.F. and T.W. were supported in part by NSFC of China under Grants No. 12125501, No. 12293060, No. 12293063, and No. 12141501, and by the National Key Research and Development Program of China under Grant No. 2020YFA0406400. B.N.L. was supported by NSAF Grant No.U2330401 and by the National Natural Science Foundation of China under Grant No.12275259.
@@ Appendix A Nucleon-nucleon Scattering
We use the spherical wall method to calculate nucleon-nucleon phase shifts on the lattice SuppNN1 ; SuppNN2 ; SuppNN3 . The fitting results of $n$-$p$ scattering for the full Hamiltonian $H$ (blue dash-dot line) and the nonperturbative Hamiltonian $H_{0}$ (black dashed line) is shown in Fig. S 1, with the center-of-mass momentum $P\leq 200$ MeV. We also show the phase shifts extracted from the Nijmegen partial wave analysis SuppNN4 , represented by the red points. For $H_{0}$, results are only fitted to $^{1}S_{0}$ and $^{3}S_{1}$ channels, so reasonable agreement with the experimental data appears in these two channels. The results for other channels are predictions based on the $S$-waves fit, and the agreement is not very satisfactory. We could in principle make an improvement by increasing the value $\Lambda^{\prime}_{\pi}$ in Eq. (13) or adding more operators in $H_{0}$, but this would introduce additional sign problems. For $H$, the fitting shows good agreement with the experiment in most channels. In several $D$-wave channels there are some deviations, which can be eliminated if operators beyond N2LO are included.
Figure S 1: $n$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ for $H$ and $H_{0}$. The red points are from Nijmegen partial wave analysis. The blue and black lines denote fitting results of the full Hamiltonian $H$ and the nonperturbative Hamiltonian $H_{0}$, respectively.
Figure S 2: $p$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ in the $^{1}S_{0}$ channel. The black points with errors are from the Nijmegen partial wave analysis. The gray line denotes the fit result.
For the isospin-breaking potentials $V_{pp}$ and $V_{nn}$, we tune the value of $C_{pp}$ to reproduce the $p$-$p$ phase shifts in the $^{1}S_{0}$ channel SuppNN4 . The fitting result is given in Fig. S 2. For $V_{nn}$, we fit it to the experimental value of the $n$-$n$ scattering length SuppNN6 and effective range SuppNN7 in the $^{1}S_{0}$ channel: